# How to be an orthodox quantum mechanic

### v2

@@ 5 Results and discussion
We can now ask how far the revealed orthodoxy diverges from the idea that it can be defined entirely by the ontological primacy of the quantum state (as advocated by Home [6]). The first element that brings this into question is the “non-ensemble $\psi$” proposition. This is a vital part of the orthodoxy (stronger even than ontological completeness) and it does not follow from $\psi$ being ontologically complete. There being no reason to preclude a fundamentally indeterminist viewpoint. If this is the case, then there are a cascade of propositions that must follow, as they all depend on “non-ensemble $\psi$” for their relevance. These are “single particles are wave-like”, “states collapse on measurement”, “measurement inherently disturbs the system”, and “measurement creates observed values”. This indicates that a single proposition reduction of the orthodoxy is unviable. The two main pillars that are empirically revealed are “measurement disturbs systems” and “states do not just describe ensembles”. Notably, the significance of consensus around ontological completeness has declined for modern works, as a result of some authors [32, 36] instead positing “statistical completeness” or openly suggesting a more general theory should be found [21].
@@ 6 Conclusion
When we consider a sample of $42$ textbooks, aimed at both graduate and undergraduate students of quantum mechanics, we find a robust orthodoxy emerging. This is powerfully illustrated by the average orthodoxity of the sources being 7373% with a standard deviation of 2424%. Given that the studied sample spans nearly a century, this provides good evidence that many aspects of the consensus have remained unchanged. In post-1980 works, the average orthodoxity is reduced to 6868% from a pre-1980 value of 8484%. In particular, there is less support for the notions that individual particles are wave-like, and that the quantum state is ontologically complete. There are also more significant reductions in support for the notions that the uncertainty principle applies to individual particles, quantum uncertainty results from measurement disturbance, particles do not have trajectories, and complementarity. The reduction in orthodoxy over time mirrors recent surveys of physicists attitudes in quantum mechanics [2]. Perhaps indicating the start of opinion fragmenting away from the orthodoxy, as there are no obvious patterns for groups of propositions being denied together.
@@
The revealed orthodoxy strongly resembles many aspects of Copenhagen-like frameworks. Particularly, there is powerful agreement (assent frequency $0.91$) that measurement inherently disturbs quantum systems, single instances of systems are individually described by a quantum state, rather than as an ensemble (assent frequency $0.96$), state collapse happens (assent frequency $0.79$), that individual particles are wave-like (assent frequency $0.88$), and that quantum states are ontologically complete (assent frequency $0.85$). These positions are further supplemented by the existence of a wave-particle duality (assent frequency $0.78$) and that measurement actually creates or alters the definite property values being measured (assent frequency $0.76$, but only $0.59$ without inferred values). More contentious, but still popular, are the notions that the uncertainty principle arises from measurement disturbance (assent frequency $0.67$), and that individual particle measurements are subject to the uncertainty principle (assent frequency $0.85$).
@@
It is clear that the orthodoxy favours the view of the quantum state as a complete description of the actual state of affairs, rather than the more instrumental view attributed to Copenhagen by authors like Stapp [4]. In this regard, the views of the orthodoxy espoused by Home [6] and Jammer [5] are vindicated. The work most embodying Stapp’s version would be Peres [38], with similarities in Mandl [39], Nielsen & Chuang [18], Hannabuss [17], and Sakurai [42]. However, it is not clear that all the logical consequences of Home’s view (that the orthodoxy is entirely characterised by ontological completeness) are believed by the physics community (and it is clear there are wave-related addenda). Thus, the orthodoxy’s characterisation requires more nuance than being distilled to a single point about the status of quantum states. It is clear that, examining all the propositions in this work, there is no room in the orthodoxy for ensemble views, entirely in opposition to the stance taken by Home. In fact, it is a strong requirement of the orthodoxy that quantum states describe single instances, rather than ensembles.
@@
It is also interesting to note that respondents in [1] characterise the Copenhagen interpretation’s most salient features as “collapse” (77%), “complementarity” (71%), “indeterminism” (46%), and “correspondence” (43%). This is quite different to the orthodoxy deduced from our sample of textbooks. Where complementarity is far less important, indeterminism is more strongly supported (non-ensemble $\psi$ + ontological completeness), and collapse is popular but not absolutely primary. This suggests that, although physicists might predominantly report subscribing to the Copenhagen interpretation, the prevailing community beliefs do not fully conform to their own characterisation of this viewpoint.

### v3

@@ 5 Results and discussion
We can now ask how far the revealed orthodoxy diverges from the idea that it can be defined entirely by the ontological primacy of the quantum state (as advocated by Home [Home1997]). The first element that brings this into question is the “non-ensemble $\psi$” proposition. This is a vital part of the orthodoxy (stronger even than ontological completeness) and it does not follow from $\psi$ being ontologically complete. There being no reason to preclude a fundamentally indeterminist viewpoint. If this is the case, then there are a cascade of propositions that must follow, as they all depend on “non-ensemble $\psi$” for their relevance. These are “single particles are wave-like”, “states collapse on measurement”, “measurement inherently disturbs the system”, and “measurement creates observed values”. This indicates that a single proposition reduction of the orthodoxy is unviable. The two main pillars that are empirically revealed are “measurement disturbs systems” and “states do not just describe ensembles”. Notably, the significance of consensus around ontological completeness has declined for modern works, as a result of some authors [phillips2013introduction, moore1998] instead positing “statistical completeness” or openly suggesting a more general theory should be found [weinberg2013].
@@ 6 Conclusion
When we consider a sample of $42$ textbooks, aimed at both graduate and undergraduate students of quantum mechanics, we find evidence of robust orthodoxy that has begun to fragment. This is powerfully illustrated by post-1980 works having an average orthodoxity of 6060%, down from a pre-1980 value of 7777%. The standard deviation almost doubles from 1616% to 2828% when moving from pre to post-1980 samples. For the full sample, the average orthodoxity of the sources is 6666% with a standard deviation of 2525%. Over time, there has been less support for the notions that individual particles are wave-like, and that the quantum state is ontologically complete. There are also more significant reductions in support for the notions that the uncertainty principle applies to individual particles, quantum uncertainty results from measurement disturbance, particles do not have trajectories, and complementarity. The reduction in orthodoxy over time mirrors recent surveys of physicists attitudes in quantum mechanics [survey2025]. Perhaps indicating the start of opinion fragmenting away from the orthodoxy, as there are no obvious patterns for groups of propositions being denied together.
@@
For the full sample of works, the revealed orthodoxy strongly resembles many aspects of Copenhagen-like frameworks. Particularly, there is powerful agreement (assent frequency $0.91$) that measurement inherently disturbs quantum systems, quantum systems have no pre-measurement values ($0.91$ assent), single instances of systems are individually described by a quantum state, rather than as an ensemble (assent frequency $0.96$), state collapse happens (assent frequency $0.79$), that individual particles are wave-like (assent frequency $0.88$), and that quantum states are ontologically complete (assent frequency $0.85$). These positions are further supplemented by the existence of a wave-particle duality (assent frequency $0.76$) and that individual particle measurements are subject to the uncertainty principle (assent frequency $0.85$). More contentious, but still popular, is the notion that the uncertainty principle arises from measurement disturbance (assent frequency $0.67$).
@@
It is clear that the orthodoxy favours the view of the quantum state as a complete description of the actual state of affairs, rather than the more instrumental view attributed to Copenhagen by authors like Stapp [stapp1972]. In this regard, the views of the orthodoxy espoused by Home [Home1997] and Jammer [Jammer1974-JAMTPO-10] are vindicated. The work most embodying Stapp’s version would be Peres [peres1993], with similarities in Mandl [mandl2013quantum], Nielsen & Chuang [nielsen2010], Hannabuss [Hannabuss:1999fa], and Sakurai [Sakurai:1167961]. However, it is not clear that all the logical consequences of Home’s view (that the orthodoxy is entirely characterised by ontological completeness) are believed by the physics community (and it is clear there are wave-related addenda). Thus, the orthodoxy’s characterisation requires more nuance than being distilled to a single point about the status of quantum states. It is clear that, examining all the propositions in this work, there is no room in the orthodoxy for ensemble views, entirely in opposition to the stance taken by Home. In fact, it is a strong requirement of the orthodoxy that quantum states describe single instances, rather than ensembles.
@@
It is also interesting to note that respondents in [sivasundaram2016surveyingattitudesphysicistsconcerning] characterise the Copenhagen interpretation’s most salient features as “collapse” (77%), “complementarity” (71%), “indeterminism” (46%), and “correspondence” (43%). This is quite different to the orthodoxy deduced from our sample of textbooks. Where complementarity is far less important, indeterminism is more strongly supported (non-ensemble $\psi$ + no pre-measurement values), and collapse is popular but not absolutely primary. This suggests that, although physicists might predominantly report subscribing to the Copenhagen interpretation, the prevailing community beliefs do not fully conform to their own characterisation of this viewpoint.